OpenStudy (anonymous):

3x + 4/x+2 - x + 3/x + 2

OpenStudy (btaylor):

since you have the same denominator, you can combine the numerators --> \[=\frac{3x + 4 -x -3}{x+2}=\frac{2x+1}{x-2}\]